OpenStudy (anonymous):

How do you find the terms of a sequence when given the formula an=5n+1 ?

OpenStudy (nincompoop):

figure out what the formal stands for

OpenStudy (nincompoop):

n = ?

OpenStudy (nincompoop):

a = ?

OpenStudy (campbell_st):

well this appears to be an aritmethic sequence \[n = 1....a_{1} = 6\] \[n = 2 ... a_{2} = 5 \times 2 + 1 = 11\] and so forth... \[n = 3... a_{3} = 5 \times 3 + 1= 16\] just keep following the pattern